# Identifying Equivalent Training Dynamics

### v2

@@ A.2 Numerical experiments
To validate that the numerically computed Koopman spectral objects, corresponding to OMD and OGD applied to specific problems, encode sufficient information to identify the conjugacy, we chose to test on the example of log barrier regularization with exponential reparameterization, as presented in [26]. In particular, we set $R=-\sum_{i=1}^{d}\log(x_{i})$, $\mathcal{K}=[0.01,1.0]^{d}$, (with $d=2$), and $q(u)=\exp(u)$. This is an example of a nonlinear conjugacy.
@@
We applied OMD and OGD on $f=\sum_{i=1}^{d}x_{i}^{4}$ (Fig. 2) and $f=\sum_{i=1}^{d}\tan(x_{i})$ (Fig. S1), for $T=100$ time steps, with a learning rate of $\eta=0.01$. We evolved 25 initial conditions, with $x(0)\in\{0.1,0.3,0.5,0.7,0.9\}\times\{0.1,0.3,0.5,0.7,0.9\}$ and $u(0)\in\{-2.30,-1.75,-1.20,-0.65,-0.10\}\times\{-2.30,-1.75,-1.20,-0.65,-0.10\}$. Using the resulting trajectories, we computed the the KMD (see Appendix A.3). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at [link to be added upon acceptance].
Figure S1: Conjugacy between online mirror descent and online gradient descent holds when finding the minimum of another function. Same as Fig. 2, but with $f(x)=\tan(x)$.
A.3 Computing Koopman mode decomposition
@@
To compute the KMD associated with optimization of OMD and OGD on $f(x)=\sum_{i=1}^{d}x_{i}^{4}$ (Fig. 2) and $f(x)=\sum_{i=1}^{d}\tan(x_{i})$ (Fig. S1), we saved the values of $x(t)$ and $u(t)$ across the $T=100$ training steps. These were concatenated into tensors $X,U\in\mathbb{R}^{2\times 100\times 25}$. Four time-delays [43, 49, 42], $d=4$, were applied, and the resulting tensors were flattened along the last dimension. This led to matrices $Z_{X},Z_{U}\in\mathbb{R}^{10\times 2375}$. We applied DMD-RRR [51] on these matrices to compute the Koopman eigenvalues.
Appendix B Fully connected neural networks
B.1 Training experiment details
@@
FCNs with only a single hidden layer are trained on MNIST, for one epoch, using SGD. Training was performed using PyTorch. All hyper-parameters used for training are presented in Table S1. All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at [link to be added upon acceptance].
@@
B.2 Computing Koopman mode decomposition
@@
As noted in Sec. 4.2, a challenge in computing the KMD associated with training FCNs is that, in order to accurately approximate the Koopman eigenvalues, multiple trajectories must be sampled. However, given that FCNs (and other DNN models) can have loss landscapes with multiple local minima, training from different random initializations can lead to different trajectories with different properties. To address this, we perform the following three steps:
@@
  • We randomly sample an initialization for the weights projecting from the inputs to the hidden layer, and the weights from the hidden layer to the outputs. We use the standard PyTorch initialization scheme, with weights being uniformly sampled in $[-\sqrt{k},\sqrt{k}]$, where $k$ is the number of input features. We denote this initialization by $W_{0}$. We then train the FCN, from $W_{0}$, for a full epoch.
@@
  • We sample another $n_{s}-1$ initializations of the FCN. Instead of randomly sampling a new set of parameters, we consider a perturbation of $W_{0}$. Namely, $W_{0}[1+\varepsilon\mathcal{N}(0,1)]$, where $\mathcal{N}(0,1)$ is a Gaussian distribution with zero mean and unit variance. The FCNs were then trained from these initializations, using the same batch order as was used to train the FCN from $W_{0}$. In our experiments, we set $n_{s}=10$ and $\varepsilon=0.001$. To investigate whether training from the perturbed initialization did indeed lead to dynamics that were in the same basin of attraction as training from $W_{0}$, we computed the ratio of their end test loss with the end test loss of the network initialized at $W_{0}$. We find that the ratio is centered around $1$ (Fig. S2), providing evidence for that the training dynamics are restricted to the basin of attraction of the same local minima.
@@
  • We repeated the steps above $n_{n}-1$ times. The weight evolutions, from each set of networks, were saved separately and the KMD was computed on each one independently. In our experiments, we set $n_{n}=25$.
@@
Figure S2: Perturbing network leads to similar loss relative to original model. To evaluate whether training the FCN from the perturbed initialization led to trajectories that lay within the same basin of attraction as the unperturbed initialization, the relative test loss (test loss for perturbed initialization divided by test loss for unperturbed initialization). For all FCN widths, we see that the distribution of relative test loss is peaked at $1$, demonstrating that the perturbed FCNs converge to a similar test loss as the unperturbed initialization. This provides evidence for the trajectories occurring in the same basin of attraction.
@@
To compute the KMD, we considered as observables the weights from the hidden layer to the output layer. This choice was made because: 1) the values of these weights determine the weight evolution of the earlier layers (due to backprop); 2) there are fewer weights from the hidden layer to the output layer, than there are from the input layer to the hidden layer. To enrich our observables, we considered time-delays [43, 49, 42]. Because we considered FCNs of differing width, using the same number of time-delays leads to matrices of different dimension. Therefore, we fixed the ratio of $d$ (number of time-delays) to $h$ (the number of units in the hidden layer). $d=64$ for FCNs with $h=5$, $d=32$ for FCNs with $h=10$, and $d=8$ for FCNs with $h=40$.
@@
We applied DMD-RRR on these time-delayed observables. When principal component analysis was performed on these observables, it was found that a lower-dimensional subspace contained a large percentage of the variance. Therefore, we considered only the top $10$ Koopman modes, using a reduced singular value decomposition (SVD).
Appendix C Convolutional neural network training phases
C.1 Training experiment details
@@
LeNet [55] and ResNet-20 [56] models were trained on MNIST and CIFAR-10 (respectively), for $20$ epochs. Training was performed using the ShrinkBench framework [69], which makes use of PyTorch. The open source code can be found here: https://github.com/JJGO/shrinkbench/tree/master. All hyper-parameters used for training are presented in Table S2. They were chosen to be the same between the two architectures to make the comparison between the two fairest, and selected the same hyper-parameters as was previously used to study LeNet [11]. Three independent seeds were trained ($n_{n}=3$), each of which was initialized from $10$ perturbed initializations ($n_{s}=10$). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at [link to be added upon acceptance].
@@
Table S2: Hyper-parameters used for CCN training in Sec. 4.3.
C.2 Computing Koopman mode decomposition
As with the FCNs (Sec. B.2), the observables used to construct the Koopman operator were time-delays of the weights going from the last hidden layer to the output layer. Eight time-delays, $d=8$, were used and the top 10 Koopman modes were considered, based on a reduced SVD. The appropriateness of this was again verified by examining the amount of variance captured by the first 10 principal components.
Appendix D Transformer grokking
D.1 Training experiment details
Single hidden layer Transformers, with four attention heads, were trained on modular arithmetic using open source code [57]: https://github.com/KindXiaoming/Omnigrok/tree/main. As noted in the repository, this itself is a modified version of previously developed code [29]. We keep all parameters the same. Therefore, we refer the interested reader to the details presented in the repository and related papers. 20 independent seeds were trained ($n_{n}=20$), each of which was initialized from 10 perturbed initializations ($n_{s}=10$). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at [link to be added upon acceptance].
@@ D.2 Computing Koopman mode decomposition
As with the FCNs (Appendix B.2) and CNNs (Appendix C.2), we use as observables time-delays of the weights from the last hidden layer to the output. $d=32$ time-delays were used. Because the number of weights was $>65000$, before performing the time-delays and flattening the tensor that stored all the weights, we performed dimensionality reduction. This was achieved by applying PCA to the flattened tensor that contained all weights (without time-delays), and then projecting the weight trajectories corresponding to the training of each perturbed initialization onto the top 10 principal components. These dimensionally reduced weights were then time-delayed and used to construct the KMD. The top 10 Koopman modes were considered, based on a reduced SVD. The appropriateness of this was again verified by examining the amount of variance captured by the first 10 principal components.

### v3

@@ A.3 Numerical experiments
To validate that the numerically computed Koopman eigenvalues, corresponding to OMD and OGD applied to specific problems, encode sufficient information to correctly identify the conjugacy, we chose to test them on the example of log barrier regularization with exponential reparameterization, as presented in Ghai et al. (2022) [26]. In particular, we set $R=-\sum_{i=1}^{d}\log(x_{i})$, $\mathcal{K}=[0.01,1.0]^{d}$, $\mathcal{K}^{\prime}=[-4.6,0.0]^{d}$, and $\mathcal{K}^{\prime\prime}=[-4/3,8/7]^{d}$ (with $d=2$), and $q(u)=\exp(u)$. This is an example of a nonlinear conjugacy.
@@
We applied OMD and OGD on $f=\sum_{i=1}^{d}x_{i}^{4}$ (Fig. S1) and $f=\sum_{i=1}^{d}\tan(x_{i})$ (Fig. 2), for $T=100$ time steps, with a learning rate of $\eta=0.01$. We evolved 25 initial conditions, with $x(0)\in\{0.1,0.3,0.5,0.7,0.9\}\times\{0.1,0.3,0.5,0.7,0.9\}$ and $u(0)\in\{-2.30,-1.75,-1.20,-0.65,-0.10\}\times\{-2.30,-1.75,-1.20,-0.65,-0.10\}$. When we used the BM (Fig. 2), we sampled 25 initial conditions, with $a(0)\in\{-16/12,-13/12,-10/12,-7/12,-4/12\}\times\{-16/12,-13/12,-10/12,-7/12,-4/12\}$ and $b(0)\in\{1/7,0.393,0.643,0.893,8/7\}\times\{1/7,0.393,0.643,0.893,8/7\}$. Note that, for simplicity, we consider each $a(0)$ and $b(0)$ only once (leaving $25$ initial conditions). Using the resulting trajectories, we computed the KMD (see Appendix A.4). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at https://github.com/william-redman/Identifying_Equivalent_Training_Dynamics.
Figure S1: Conjugacy between online mirror descent and online gradient descent is identifiable from Koopman spectra. (A) Comparing example trajectories of variables optimized via OMD ($x_{1},x_{2}$) and OGD ($u_{1},u_{2}$), the existence of a conjugacy is not obvious. (B) Similarly, the existence of a conjugacy is not apparent when looking at the loss incurred by using OMD and OGD. (C) Comparing the Koopman eigenvalues associated with optimizing using OMD and OGD correctly identifies the existence of a conjugacy. The function optimized is in all subfigures is $f(x)=\sum x^{4}$. Note that the BM was not used in this figure as $f(x)$ is symmetric around its minimum, which does not satisfy the assumption that $f(a)<0$.
A.4 Computing the Koopman mode decomposition
@@
To compute the KMD associated with optimization using OMD and OGD on $f(x)=\sum_{i=1}^{d}\tan(x_{i})$ (Fig. 2) and $f(x)=\sum_{i=1}^{d}x_{i}^{4}$ (Fig. S1), we saved the values of $x(t)$ and $u(t)$ across the $T=100$ training steps. These were concatenated into tensors $X,U\in\mathbb{R}^{2\times 100\times 25}$. Four time-delays [44, 45, 51] were applied, and the resulting tensors were flattened along the last dimension. This led to matrices $Z_{X},Z_{U}\in\mathbb{R}^{10\times 2375}$. We applied DMD-RRR [53] on these matrices to compute the Koopman eigenvalues. The same approach was used to compute the KMD associated with optimization via BM.
Appendix B Randomized shuffle control
While the Wasserstein distance gives a natural way to quantify how similar the Koopman spectra associated with two dynamical systems are, it remains an open question on how to best interpret the magnitude of the computed distance. In particular, if $\Lambda^{(1)}=\{\lambda_{1}^{(1)},...,\lambda_{N}^{(1)}\}$ and $\Lambda^{(2)}=\{\lambda_{1}^{(2)},...,\lambda_{N}^{(2)}\}$, what value of the Wasserstein distance $\omega=W_{2}\left(\Lambda^{(1)},\Lambda^{(2)}\right)$ is “sufficiently small” enough that we can confidently conclude the two dynamical systems are topologically conjugate?
To help increase the transparency and interpretability of our results, we develop a randomized shuffle control to act as a baseline to assess how significantly distinct $\Lambda^{(1)}$ and $\Lambda^{(2)}$ are. In particular, by comparing $\omega$ to a distribution of Wasserstein distances computed from randomly shuffling $\Lambda^{(1)}$ and $\Lambda^{(2)}$, we may estimate whether the true Koopman eigenvalues are more or less “distant” than would be expected by chance.
One challenge that emerges in creating a good randomized shuffle control is the fact that the Koopman eigenvalues correspond to dynamical properties of the underlying system (e.g., decay rates, oscillations, fixed points). Therefore, if we take the naive approach and randomly assign half of $\Lambda=\{\Lambda^{(1)},\Lambda^{(2)}\}$ to $\Lambda^{\prime(1)}$ and the other half to $\Lambda^{\prime(2)}$, then $\Lambda^{\prime(1)}$ and $\Lambda^{\prime(2)}$ could be “non-natural”, having only slow or fast decays modes, no fixed points, etc. In such a case, we expect $\omega<\omega^{\prime}=W_{2}\left(\Lambda^{(1)},\Lambda^{(2)}\right)$ in general. This makes the naive approach a poor baseline.
To attempt to construct a stronger and more informative baseline, we randomly assign eigenvalues to $\Lambda^{\prime(1)}$ and $\Lambda^{\prime(2)}$ in the following manner. First, by computing $W_{2}\left(\Lambda^{(1)},\Lambda^{(2)}\right)$ we identify the “assignment” $\sigma:\{1,...,N\}\rightarrow\{1,...,N\}$ that maps the order of $\Lambda^{(2)}$ to be as close to $\Lambda^{(1)}$ as possible. In particular, $\sigma$ is defined as
$\min_{\sigma}\sum_{i=1}^{N}||\lambda^{(1)}_{i}-\lambda^{(1)}_{\sigma(i)}||_{2}.$ (8)
To generate the shuffled eigenvalues, we then place $\lambda_{i}^{(1)}$ in $\Lambda^{\prime(1)}$ with $50\%$ probability and in $\Lambda^{\prime(2)}$ with $50\%$ probability. The eigenvalue in $\Lambda^{(2)}$ that is closest to $\lambda_{i}^{(1)}$ via the assignment $\sigma$ [i.e., $\lambda_{\sigma(i)}^{(2)}$] is then placed in whichever $\Lambda^{\prime(1)}$ or $\Lambda^{\prime(2)}$ that $\lambda_{i}^{(1)}$ is not. Performing $n_{\text{shuff}}=100$ shuffles, we get a distribution of Wasserstein distances between the shuffled eigenvalues, $\{\omega^{\prime}_{1},...,\omega^{\prime}_{n_{\text{shuff}}}\}$. We report the number of shuffles that have a Wasserstein distance greater than or equal to $\omega$. In the case where the two sets of Koopman eigenvalues are highly distinct, we expect there to be few shuffles that satisfy $\omega^{\prime}\geq\omega$. In contrast, when the Koopman eigenvalues are highly overlapping, we expect there to be more shuffles that satisfy $\omega^{\prime}\geq\omega$. For this reason, we report the percent of shuffles with $\omega^{\prime}\geq\omega$ in the main text.
This approach to defining the randomized shuffle control has two useful properties. First, $\Lambda^{\prime(1)}$ and $\Lambda^{\prime(2)}$ have as similar magnitudes of eigenvalues to $\Lambda^{(1)}$ and $\Lambda^{(2)}$ as possible. And second, if $\lambda^{(1)}_{i}$ and $\lambda^{(1)}_{i+1}$ are complex conjugate pair of eigenvalues, whose closest matches are $\lambda^{(2)}_{\sigma(i)}$ and $\lambda^{(2)}_{\sigma(i+1)}$ that are real only, then the complex conjugate pair can be “split” when creating $\Lambda^{\prime(1)}$ and $\Lambda^{\prime(2)}$. This can lead to $\omega^{\prime}<\omega$. Thus, there is a “penalty” when the number of complex conjugate eigenvalues does not match between $\Lambda^{(1)}$ and $\Lambda^{(2)}$.
We note that this randomized shuffle control is not a perfect construct and represents one possible way of generating significance. We hope that future work will enable rigorous comparison between Koopman eigenvalues by developing computation schemes for estimating how much two dynamical systems diverge given the differences in their approximate spectra.
Appendix C Fully connected neural networks
C.1 Training experiment details
@@
FCNs with only a single hidden layer are trained on MNIST, for one epoch, using SGD. Training was performed using PyTorch. All hyper-parameters used for training are presented in Table S1. All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at https://github.com/william-redman/Identifying_Equivalent_Training_Dynamics.
@@
C.2 Computing the Koopman mode decomposition
@@
A challenge in computing the KMD associated with training FCNs is that, in order to accurately approximate the Koopman eigenvalues, multiple trajectories must be sampled. However, given that FCNs (and other DNN models) can have loss landscapes with multiple local minima, training from different random initializations can lead to different trajectories with different dynamical properties. To address this, we perform the following three steps:
@@
  • We randomly sample an initialization for the network input and output weights. We use the standard PyTorch initialization scheme, with weights being uniformly sampled in $[-\sqrt{k},\sqrt{k}]$, where $k$ is the number of input features. We denote this initialization by $W_{0}$. We then train the FCN, from $W_{0}$, for a full epoch.
@@
  • We sample another $n_{s}-1$ initializations of the FCN, for $n_{s}\in\mathbb{N}$. Instead of randomly sampling a new set of parameters, we consider a perturbation of $W_{0}$. Namely, $W_{0}[1+\varepsilon\mathcal{N}(0,1)]$, where $\mathcal{N}(0,1)$ is a Gaussian distribution with zero mean and unit variance. The FCNs were then trained from these initializations, using the same batch order as the one used to train the FCN from $W_{0}$. In our experiments, we set $n_{s}=10$ and $\varepsilon=0.001$. To investigate whether training from the perturbed initialization did indeed lead to dynamics that were in the same basin of attraction as training from $W_{0}$, we computed the ratio of their end test loss with the end test loss of the network initialized at $W_{0}$. We find that the ratio is centered around $1$ (Fig. S2), providing evidence for that the training dynamics are restricted to the basin of attraction of the same local minimum.
@@
  • We repeated the steps above $n_{n}-1$ times, for $n_{n}\in\mathbb{N}$. The weight evolutions, from each set of networks, were saved separately and the KMD was computed on each one independently. In our experiments, we set $n_{n}=25$.
@@
Figure S2: Perturbing network weights leads to similar loss relative to original model. To evaluate whether training the FCN from the perturbed initialization led to trajectories that lay within the same basin of attraction as the unperturbed initialization, we compute the relative test loss (test loss for perturbed initialization divided by test loss for unperturbed initialization). For all FCN widths, we see that the distribution of relative test loss is peaked at $1$, demonstrating that the perturbed FCNs converge to a similar test loss as the unperturbed initialization. This provides evidence that the trajectories lie in the same local minimum basin.
@@
To compute the KMD, we considered as observables the weights from the hidden layer to the output layer. This choice was made because: 1) the values of these weights determine the weight evolution of the earlier layers (due to backprop); 2) there are fewer weights from the hidden layer to the output layer, than there are from the input layer to the hidden layer, enabling our approach to be more computationally tractable. To enrich our observables, we considered time-delays [44, 45, 51]. Because we considered FCNs of differing width, using the same number of time-delays leads to matrices of different dimension. Therefore, we fixed the ratio of $d$ (number of time-delays) to $h$ (the number of units in the hidden layer), setting $d=64$ for FCNs with $h=5$, $d=32$ for FCNs with $h=10$, and $d=8$ for FCNs with $h=40$.
@@
We applied DMD-RRR [53] on these time-delayed observables. When principal component analysis (PCA) was performed on these observables, it was found that a lower-dimensional subspace contained a large percentage of the variance. This is in-line with previous work finding the weights of DNNs traverse low dimensional subspaces [3]. Therefore, we considered only the top $10$ Koopman modes, using a reduced singular value decomposition (SVD).
C.3 GeLU FCNs
To understand how robust our conclusion that shallow and wide FCNs have non-conjugate training dynamics is, we perform the same Koopman-based analysis on FCNs with GeLU activation functions [59]. Given the similarity between ReLU and GeLU, we expect that they should have similar training dynamics behavior. Computing the Koopman eigenvalues and comparing across FCNs of varying widths, we find very similar results (compare Fig. S3 with Fig. 3). This demonstrates that our Koopman based framework is robust to (minor) hyper-parameter differences. A full examination of how other choices of activation function (especially those that introduce “squashing” – e.g., sigmoid) impact the training dynamics would be an interesting future direction to pursue.
Figure S3: Narrow and wide fully connected neural networks have non-conjugate training dynamics when using GeLU activation functions. Same as Fig. 3D–F in the main text, but for FCNs using GeLU activation functions. Error bars are $\pm$ standard deviation across $10$ independently trained FCNs.
C.4 Conjugate training dynamics across random initializations of FCNs
Motivated by recent work arguing that different random initializations of DNNs converge to solutions with no loss barrier in the linear interpolation between them [16], when taking into account the inherent permutation symmetry of the parameters, we asked whether our Koopman-based framework identified conjugate training dynamics for different initialized FCNs. We examine how similar the Koopman spectra for all 25 independently trained FCNs of width $h=40$ are. We chose this because Entezari et al. (2022) [16] proved their result for sufficiently wide FCNs, but sufficiently narrow FCNs exhibited a loss barrier along the linear interpolation, when permutation symmetry was not taken into account [16]. Our $h=40$ FCNs strike a balance between these two points, making it a good point for analysis.
We find examples of different random initializations with nearly identical Koopman spectra (Fig. S4A). Across all possible pairs of the 25 independently trained FCNs, we find that the Wasserstein distance between Koopman spectra is centered around values similar to what was seen when comparing $h=10$ and $h=40$ FCNs. These results support the hypothesis that, when taking into account permutation symmetry (which computation of the Koopman eigenvalues naturally does), at least some random initializations have conjugate training dynamics.
Figure S4: Conjugate training dynamics across random initializations of FCNs. (A) Example Koopman spectra associated with training FCNs, $h=40$, from two different random seeds. (B) Distribution of Wasserstein distance between the Koopman spectra of all possible pairs of 25 independently trained FCNs.
Appendix D Convolutional neural network training phases
D.1 Training experiment details
@@
LeNet [60] and ResNet-20 [61] models were trained on MNIST and CIFAR-10 (respectively), for $20$ epochs. Training was performed using the ShrinkBench framework [74], which makes use of PyTorch. The open source code can be found here: https://github.com/JJGO/shrinkbench/tree/master. All hyper-parameters used for training are presented in Table S2. They were chosen to be the same between the two architectures to make the comparison between the two fair and were selected to match the hyper-parameters that were previously used to study LeNet’s training dynamics [11]. Three independent seeds were trained ($n_{n}=3$), each of which was initialized from $10$ perturbed initializations ($n_{s}=10$). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at https://github.com/william-redman/Identifying_Equivalent_Training_Dynamics.
@@
Table S2: Hyper-parameters used for CNN training in Sec. 4.3.
@@ D.2 Computing Koopman mode decomposition
As with the FCNs (Sec. C.2), the observables used to construct the Koopman operator were time-delays of the weights going from the last hidden layer to the output layer. Eight time-delays, $d=8$, were used and the top 10 Koopman modes were considered, based on a reduced SVD. The appropriateness of this was again verified by examining the amount of variance captured by the first 10 principal components.
D.3 Evaluating CNN training phases at a coarser timescale
In addition to studying the earliest part of CNN training, which was the focus of previous work [27], we examined how CNN training dynamics evolved across the first 20 epochs of training. To reduce the computational cost associated with computing and comparing Koopman spectra, we computed the Koopman mode decomposition across all training iterations within a single epoch.
At this coarser timescale, we find that LeNet, trained on MNIST, and ResNet-20, trained on CIFAR-10, exhibit very similar evolutions in training dynamics (Fig. S5). Indeed, both see a continual reduction in Wasserstein distance between neighboring epochs as training time goes on. Interestingly, in both cases, there is a growing block diagonal of dark blue, that becomes especially strong at training epoch 13. This supports the general similarity in the evolution of training dynamics between different CNN architectures, as was previously observed [27].
Figure S5: Koopman-based framework enables identification of transitions in dynamics across the training of LeNet and ResNet-20. (A)–(B) Same as Fig. 4A–B, but for dynamics computed over individual epochs.
Appendix E Transformer grokking
E.1 Training experiment details
Single hidden layer Transformers, with four attention heads, were trained on modular arithmetic using open source code [62]: https://github.com/KindXiaoming/Omnigrok/tree/main. As noted in the repository, this is a modified version of previously developed code [29]. We keep all hyper-parameters the same. Therefore, we refer the interested reader to the details presented in the repository and related papers. 20 independent seeds were trained ($n_{n}=20$), each of which was initialized from 10 perturbed initializations ($n_{s}=10$). All experiments were run on a MacBook Air with an Apple M1 chip, 1 CPU, and no GPUs. Code implementing our experiments can be found at https://github.com/william-redman/Identifying_Equivalent_Training_Dynamics.
E.2 Computing Koopman mode decomposition
As with the FCNs (Appendix C.2) and CNNs (Appendix D.2), we use as observables time-delays of the weights from the last hidden layer to the output. $d=32$ time-delays were used. Because the number of weights was $>65000$, before performing the time-delays and flattening the tensor that stored all the weights, we performed dimensionality reduction. This was achieved by applying PCA to the flattened tensor that contained all weights (without time-delays), and then projecting the weight trajectories corresponding to the training of each perturbed initialization onto the top 10 principal components. These dimensionally reduced weights were then time-delayed and used to construct the KMD. The top 10 Koopman modes were considered, based on a reduced SVD. The appropriateness of this was again verified by examining the amount of variance captured by the first 10 principal components.